OpenStudy (anonymous):

Suppose the temperature increases 8º to -7ºF. What was the starting temperature?

OpenStudy (texaschic101):

starting temperature = x x + 8 = -7 --- subtract 8 from both sides x = -7 - 8 x = - 15 check.. -15 + 8 = -7 -7 = -7 (correct) starting temperature is -15